OpenStudy (fvf1977):

Find the value of each of the other 5 trig functions for an angle 'X', w/out finding X, given: cos(x)=-sq. rt.5/3 & cot(x)>0

OpenStudy (anonymous):



OpenStudy (anonymous):

fins the missing side via pythagoras then you can find any trig ratio you need

OpenStudy (fvf1977):

i put sinx=2/3 and it was marked wrong, any ideas?

OpenStudy (freckles):

if cot is + and cos is +, then sin is ?

OpenStudy (fvf1977):

positive?

OpenStudy (freckles):

\[\cot(x)=\frac{\cos(x)}{\sin(x)}\] cos(x)/sin(x) is positive if both cos(c) and sin(x) are positive or cos(x)/sin(x) is positive if both cos(X) and sin(x) are negative

OpenStudy (freckles):

well okay and if cot is + and cos is -, then sin is?

OpenStudy (fvf1977):

negative

OpenStudy (freckles):

right you have here in your problem cot is positive and cos is negative so your sin value will be negative as well

OpenStudy (fvf1977):

if you draw it on a graph, will it be in quadrant 2? if it is wouldnt the adjacent side be negative?

OpenStudy (freckles):

you will be in quadrant 3 because both sin and cos are negative

OpenStudy (freckles):